Roger is travelling with a speed of 0.800c relative to Mia. Roger travels for 50.0 s as measured on his watch.

 Calculate the elapsed time on Mia’s watch during this motion.

c = 3.0 ×

108

1
Expert's answer
2022-04-22T16:43:26-0400

"t=\frac{t_0}{\sqrt{1-\frac{v^2}{c^2}}}=83~s."